After leaving Moose Jae, Hanson drove 100 kilometers at x km/hr and then decreased his speed by 20 km/hr and drove another 200 additional kilometers. Write a rational expression for his total traveling time.
Evaluate the expression if x=100

Answer by AnlytcPhil(1116)

>>...After leaving Moose Jae, Hanson drove 100 kilometers at 
x km/hr and then decreased his speed by 20 km/hr and drove
another 200 additional kilometers. Write a rational expression
for his total traveling time.<<
Use TIME = DISTANCE/SPEED
>>...drove 100 kilometers at x km/hr...<<
Translation: TIME to travel the 100 kilometers = DISTANCE/SPEED
or 100/x

>>...decreased his speed by 20 km/hr and drove another 200
additional kilometers...<<
Translation: He decreased his speed from x km/hr to (x-20) km/hr
So the TIME for the 200 additional kilometers = DISTANCE/SPEED
or 200/(x-20)
So:
Total TIME = TIME to travel the 100 km + TIME to travel the 200 km
TOTAL TIME = 100/x + 200/(x-20)
>>...Evaluate the expression if x=100...<<
TOTAL TIME = 100/100 + 200/(100-20)
= 1 + 200/80
= 1 + 2.5
= 3.5 hours.
Edwin J